OpenStudy (anonymous):

Solve: 2(7 – x) = 3(x + 3)

OpenStudy (anonymous):

14-x=3x+9?

OpenStudy (kidrah69):

yep

OpenStudy (kidrah69):

now combine like terms

OpenStudy (anonymous):

Expand 14−2x=3x+9 Add 2x to both sides 14=3x+9+2x Simplify 3x+9+2x to 5x+9 14=5x+9 Subtract 9 from both sides 14−9=5x Simplify 14−9 to 5 5=5x Divide both sides by 5 1=x Switch sides x=1

OpenStudy (anonymous):

i think thats how you do it?

OpenStudy (kidrah69):

yea thats right good job :)

OpenStudy (anonymous):

:)